OpenStudy (amenah8):

The length of a rectangle is 3 inches more than twice the width. Find the length and the width if the perimeter is 98 inches.

OpenStudy (campbell_st):

let the width be w then the length is 2w + 3 so using P = 2w + 2l you get 98 = 2w + 2(2w + 3) or 98 = 6w + 6 now just solve for w

OpenStudy (amenah8):

Thank you. Does w equal 92/6?

OpenStudy (campbell_st):

no you get 92 = 6w divide both sides by 6 that will be the width... the length is 2w + 3

OpenStudy (amenah8):

I got 15 and 1/3 (mixed number) as my answer for what w will be. My length was 33 and 1/3 (mixed number). I added the length and width together, but I didn't get 98. Where did I go wrong?

OpenStudy (amenah8):

Campbell, are you still available ø help me with my problem, above? Thank you.

OpenStudy (campbell_st):

thats correct w = 15 1/3 length = 2w + 3 so length = 2*15 1/3 + 3 = 33 2/3 hope it helps

OpenStudy (amenah8):

The length and width are supposed to add up to 98, the perimeter, but the length (33 1/3) and the width (15 1/3) don't add up to 98.

OpenStudy (amenah8):

Are you still there?

OpenStudy (campbell_st):

width is 15 1/3 and length 2w + 3 = 30 2/3 + 3 = 33 2/3